The sum of two numbers is 16. Twice the smaller # is 2 more than the larger #.
Find them. ??Can someone please help me.

Found 2 solutions by scott8148, Superownage:
Answer by scott8148(6628)

let x=smaller

"Twice the smaller # is 2 more than the larger #" __ 2x-2=larger

"The sum of two numbers is 16" __ x+2x-2=16 __ 3x=18

Answer by Superownage(7)

Hello, in simple terms x+y=16 and 2y = x +2 so x=10 and y=6 (y is the smaller digit, x is the larger)